OpenStudy (anonymous):

there are 52 blades of grass in 2 in^2 of lawn. there are 580 blades of grass in 8 in^2 of the same lawn. which equation models the y blades of grass found in x in^2 of lawn?

OpenStudy (anonymous):

....y=-88x-124 ...y=-88x+124 ....y=88x-124 ...y=88x+24

OpenStudy (anonymous):

.y=88x-124